find two consecutive positive integers such that the square of the second integer added to four times the first is equal to 41
Answer by Alan3354(69443)

find two consecutive positive integers such that the square of the second integer added to four times the first is equal to 41
-------------
n^2 + 4(n-1) = 41
n^2 + 4n - 45 = 0
(n-5)*(n+9) = 0
n = 5